OpenStudy (anonymous):

Find the oblique asymptote for the graph of y=x^3/x^(2)-x, i got x+1 am i right ??

OpenStudy (accessdenied):

what was your work to get that answer?

OpenStudy (anonymous):

i used long divison

OpenStudy (accessdenied):

so you got... x + 1 + r(x)/(x^2 - x) from long division? and then selected y = x+1 as your oblique asymptote

OpenStudy (anonymous):

yes and my resulatant was x+0

OpenStudy (accessdenied):

right. so by definition, as x approaches larger and larger numbers, our remainder is x/(x^2 - x) will simply disappear (denominator >> numerator, 1/big number --> 0). I think your answer looks good. :)

OpenStudy (accessdenied):

it helps to imagine the behavior for very very large numbers: y = x + 1 + 1/x-1 anything with larger degree in denominator than numerator approaches 0 for the same reasoning as above then y = x + 1 is the end behavior.

OpenStudy (anonymous):

what does this mean ?? What combinations of vertical, horizontal, and oblique asymptotes are not possible for a rational function?

OpenStudy (accessdenied):

what are the conditions for each type of asymptote?

OpenStudy (anonymous):

what do u mean ? like ?

OpenStudy (accessdenied):

For a vertical asymptote, you need a factor in the denominator of x that allows you to get 0 like f(x) = (x + 1)/(x - 1) x=1 is a vertical asymptote

OpenStudy (accessdenied):

what would tell you that you have a horizontal asympttote? or oblique asymptote?

OpenStudy (accessdenied):

and most relevant to your question.. can you have both based on what is required?

OpenStudy (anonymous):

confused :?

OpenStudy (anonymous):

wait i looked at my notes here what i says "the degree of the numerator P(x) must be one more than the degree of the denominator "

OpenStudy (accessdenied):

that sounds good for oblique asymptotes. what about horizontal? :)

OpenStudy (anonymous):

not sure about that

OpenStudy (accessdenied):

my best hint is that it also has to do with degrees of the numerator / denominator.

OpenStudy (accessdenied):

like, would y = 1/(x - 1) have a horizontal asymptote? y = x/(x - 1)? y = x^2 / (x - 1) ?

OpenStudy (anonymous):

not really

OpenStudy (accessdenied):

again, for big values of x: y = 1/(x-1) goes to y = 1/biiiig number, so it goes to 0. thus, horizontal asymptote is at y=0. for y = x/ (x-1), you have an equally large number for numerator and denominator (the 1 is unimportant because it is so comparatively small). horizontal asymptote is at y = 1/1 = 1 but y = x^2 / (x-1) is the same problem we just had (removing the x in numerator/denominator). it had an oblique asymptote, right?

OpenStudy (anonymous):

yea

OpenStudy (accessdenied):

the first two are degree(numerator) <= degree(denominator), and they had horizontal asymptotes. the last was degree(numerator) > degree(denominator), so that we had an oblique asymptote.

OpenStudy (accessdenied):

so back to this question: What combinations of vertical, horizontal, and oblique asymptotes are not possible for a rational function? can the degree of a numerator be both greater than - and less than/equal to the degree of the denominator to give us both horizontal and oblique asymptotes?

OpenStudy (anonymous):

its can't be less than but more than for sure

OpenStudy (accessdenied):

In a slightly more math-y perspective: \( \deg (\text{numerator} ) \le \deg ( \text{denominator} ) \) => Horizontal Asymptote \( \deg ( \text{numerator) } > \deg ( \text{denominator} ) \) ==> Oblique Asymptote If we know both of these values, can both statements be true simultaneously? (a yes/no question) it is more or less like asking: can you look left /and/ right at the same time? Hope this makes more sense like this.

OpenStudy (accessdenied):

or we just call the degree of numerator N, and degree of denominator D: \( D \le N \) and \( D > N \) these simply have nothing in common, and i invite you to try any two numbers (N,D) to see if you can make both true.